1. 74.8 g of copper at 143.20C is immersed in 165 mL of glycerin (C3H8O3, dglycerin = 1.26g / mL) at 24.80C in an insulated container. The final temperature is measured at 31.10C. Accordingly, what is the heat capacity of glycerine in the unit of J. mol-1.0C-1? (c copper = 0.385 J. mol-1.0C-1)

2. Methane, the main component of natural gas, burns as follows. Accordingly, how many kJ heat is released from the full combustion of 1.63x103L methane at 18.60C and 768mmHg pressure? Also, considering that this reaction occurs in 100% efficiency, how many liters of water can increase the temperature from 8.70C to 60.00C if used in heating?

3. Show how the enthalpy changes of the reactions below occur in accordance with the Hess Law, which formation steps occur, and what the
1
Expert's answer
2020-05-06T13:59:06-0400

1. 

Cu:

M = 63.55

n = 74.8 / 63.55 = 1,177 mol

Specific molar heat capacity - 24.5! (https://hyperphysics.phy-astr.gsu.edu/hbase/Tables/sphtt.html)

Q = ncΔT

Q1 = 1,177 * 24.5 * (143.2 - 31.1)

Q1 = 3232.6 J

Glycerin:

165 mL * 1.26 g = 207,9 g

M = 92.09 g/mol

n = 207.9 / 92.09 = 2.258 mol

Q2 = Q1 = 50,8 J = 2.258 * c * (31.1 - 24.8)

c = 227.3 J/mol*C


2.

Methane:

PV = nRT

n = PV/RT

R = 8.314

P = 102 391.579 pascals

T = 18.6 + 273.15 = 291,75 K

V = 1.63 m3

n = 68.8 mol

Enthalpy of combustion ΔcHo−882.0 kJ/mol

Q = 882.0 * 68.8 = 60681.6 kJ


Q = cmΔT

4.186 joule/gram - heat capacity, water

m = Q / cΔT

m = 60681.6 * 1000 / 4.186 * (60.0-8.7)

m = 282.6 kg equal to 282.6 L


3.

Hess's Law of Constant Heat Summation (or just Hess's Law) states that regardless of the multiple stages or steps of a reaction, the total enthalpy change for the reaction is the sum of all changes.